OpenStudy (anonymous):

Math Analysis: (csc x - cot x)(sec x +1)

OpenStudy (anonymous):

i did but at the back of my book it says that the answer is tan x :\

hero (hero):

Well, that just means to change everything back into sine and cosine and tangent, and then simplify

OpenStudy (anonymous):

could you do the steps for me please? :| I tried so many times and I still didn't get the answer from the back of the book..

OpenStudy (anonymous):

nvm i figured it out myself but thx for the help anyways :p

OpenStudy (phi):

I always change sec to 1/cos csc to 1/sin cot to cos/sin and hope for the best! in this case, (csc -cot) = (1/sin - cos/sin) = (1-cos)/sin (sec+1)= (1/cos +1) = (cos+1)/cos when you multiply them you get (1-cos^2)/(sin cos) you should know 1-cos^2 = sin^2 so it looks like you get sin^2/(sin cos) = sin/cos = tan

OpenStudy (anonymous):

thx but I already solved it xP you came to late.. :p

OpenStudy (phi):

that was for anyone who wondered how to do it. You could have posted your answer, 'cause some people might want to know how to do it.

OpenStudy (anonymous):

oh, well thx haha :) I'll remember that next time xP Thx for your time and effort~ ^^